A certain type of machine requires repairs once every two years on average.
What is the probability that such a machine will work at least two years without requiring repair?
A company has four such machines which operate independently. What is the probability that at least one of them will not need repair in the next 2 years?

Answer by edjones(8007)

1/2 the probability that such a machine will work at least two years without requiring repair.
(1/2)^4=1/16 probability all will need repair in the next 2 years.
1 - 1/16 = 15/16 the probability that at least one of them will not need repair in the next 2 years.
.
Ed